OpenStudy (anonymous):

use substitution y=2x+3 y=3x+1

OpenStudy (anonymous):

let y=2x+3 be equation 1. Okay? Now, substition means, that the value of y in equation 1 is same for equation 2?

OpenStudy (anonymous):

@Jessii I believe I have you on this?

OpenStudy (anonymous):

yeah. would you multiply the 2 and 3 with a number that they both equal to? like 3 * 2=6 and 2*3=6 but make one of them negative

OpenStudy (anonymous):

No no, we're doing it by substitution. Don't jump steps here. Trust me. =) A yes no at my first question.

OpenStudy (anonymous):

y=2x+3 is the value of y which is also valid for y=3x+1 Basically these equations represent lines. The solution will be a point where these lines would interesct.

OpenStudy (anonymous):

intersect*

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So, since y is same for both, I can say that 2x+3=3x+1 Right?

OpenStudy (anonymous):

The idea is that two lines intersect at a point, the point is same for both the lines.

OpenStudy (anonymous):

so, i'd subtract 2x from 3x then it would be 1x=3?

OpenStudy (anonymous):

or would i add 2x and 3x 5x=3

OpenStudy (anonymous):

No. See, let's take that step by step too. Subtract 2x from both sides of the equation, that is:- 3=x+1 Subract 1 from both sides of equation that is:- 2=x

OpenStudy (anonymous):

Putting this value of x in both the equations would give you a value of y. I hope this is clear.